Acetylene gas, C2H2, can be produced by reacting solid calcium carbide, CaC2, with water. The products are acetylene and calcium hydroxide, Ca(OH)2. How many grams of calcium carbide is needed to produce 1.42 liters of acetylene collected at 25°C and 715 torr of pressure. (At 25°C the vapor pressure of water is 23.8 torr) The equation is: CaH2 + 2H2O —> C2H2 + Ca(OH)2


1
Expert's answer
2021-04-27T07:54:50-0400

The formula mass of CaC2 is (40.08 (Ca) + 2 × 12.01 (C)) g mol-1 = 64.1 g mol-1

The amount of CaC2 in 1.000 g is therefore, n = mM1420 /64.1

= 22.153 mol.

P= nRt/V

273 + 25 = 298

n =0.8206 × 298/1.4

= 174.67/715

= 0.24m

22.154 + 0.24

= 22.394 × 64.1

= 1435.4554g